OpenStudy (anonymous):

Fantastic Furniture builds sofas. It costs $2,725 to build 9 sofas and $3,850 to build 14 sofas. Which equation represents the cost, C(x), as a linear function of the number of sofas, x?

OpenStudy (anonymous):

a. C(x) = 225x - 700 b. C(x) = 225x + 700 c. C(x) = 700x - 225 d. C(x) = 700x + 225

OpenStudy (anonymous):

We have two points on our line: (9, 2725) and (14, 3850). Using point-slope form: \[ \Large m = {(y_1 – y_0) \over (x_1 – x_0)} \\ \ \\ \Large m = {(3850 – 2725) \over (14 – 9)} \] So, \(m = 245\). Pluggin' it in to our slope intercept form: y = C(x) = mx + b = 245x + b So next up, we find b by plugging in the values for x and y from any point we choose. Let's choose (9, 2725): y = 245x + b (2925) = 245(9) + b b = 720 Pluggin' that back into our equation: y = C(x) = mx + b = 245x + b = 245x + 720 C(x) = 245x + 720

OpenStudy (anonymous):

But that's not an option ...

OpenStudy (anonymous):

Whoops, I messed up on my calculation for m. m=225. I'll leave the rest up to you to figure out.

OpenStudy (anonymous):

you need the line through the points \((9, 2725)\) and \((14,3580)\) calculation is annoying, just ask http://www.wolframalpha.com/input/?i=line+%289%2C2725%29%2C%2814%2C3850%29 looks like the answer is B